OpenStudy (anonymous):

the net for a cylinder is shown at the right. If the radius is 6cm and the height is 15cm find the following for the cylinder Lateral Area & Total Area?

OpenStudy (anonymous):

Total Area=2pi(6^2)+(2*6)(15)

OpenStudy (anonymous):

total area =2*pi*r(r+h) =2*3.14*6*21 =791 cm^2 lateral =2*3.14*6*15 =565.2cm^2